OpenStudy (anonymous):

Imagine that a machine enables you to do 100 joules of work with a force input of 20 newtons. Now imagine that a new machine is invented that lets you do the same amount of work with a force input of just 10 newtons.The NEW machine makes you how many times stronger than the old one?

OpenStudy (anonymous):

assume that the distance is 1 m then the output force is 100 N Mechanical advantage 1 = Fout/Fin= 100/20 = 5 Mechanical advantage 2 = 100/10= 10 now divide MA NEW / MA OLD = 10/5 = 2. (we divided the so the assumed distance wouldn't change the ratio you divided numerator and denominator over the same number) so the NEW is twice stronger than the old one.